# ECAMP: Entity-centered Context-aware Medical Vision Language Pre-training

### 20 Dec 2023 v1

@@ 5.3 Ablation Study and Analysis
Advantages of context distillation The first row in Table 4 is trained by basic MIM and MLM. When we use the distilled easy-comprehensible report, we observe a performance improvement of 3% on COVIDx and 0.5% on ChestX-ray14, validating the significant role of entity-aware context distillation.
Effectiveness of descriptors masking After masking the critical descriptive words preceding the entities, the accuracy is slightly improved (shown in Table 4). Accordingly, it is confirmed that descriptors masking can better leverage the supervised signals in medical reports.
Sensitive analysis of the re-balanced factor We evaluate ECAMP using $\lambda_{neg}$ corresponding to different imbalance ratios between positive and negative expressions. As shown in Table. 5, setting $\lambda_{neg}$ to $0.05$ according to the imbalance ratio (20:1) leads to best performance.
Performance of context-guided SR As listed in Table 4, employing context-guided SR provides performance gains on ChestX-ray from 80.7% to 81.2%, which validates its effectiveness in learning informative features by generating detailed pathology vision context.
Contributions of MSCF We exclusively integrate the global vision feature $f_{v}^{g}$ and local aggregated feature $f_{a}^{l}$ to predict the masked words respectively. The results in Table 4 validate that, only fusing $f_{v}^{g}$ decreases the performances on low-level vision tasks, while the classification performance relies on well-trained $f_{a}^{l}$. It is confirmed that MSCF successfully contributes to concurrently achieving SOTA performance on multi-scale downstream tasks.
Figure 4: Attention maps learned by local aggregated features in MSCF, which successfully localizes the pathology region of atelectasis.
Representation analysis As the attention map of the local vision feature and the corresponding distilled reports visualized in Fig. 4, significant correspondence around correct pathology regions (marked as red boxes) is captured. This supports the MSCF mechanism of learning discriminative fine-grained representations, which is pivotal for improving the performance in the local downstream tasks.
@@ 6 Impact, Limitation, and Conclusion
Appendix A Statistics of descriptive words
As illustrated in Fig. S1, we present the statistics of our defined descriptive words across 227,835 reports. Specifically, we focus on descriptors with a frequency greater than 10,000 among the total of 31,597 terms. Notably, only ”may be” and ”there is” imply uncertainty about the existence of the entity, such as a disease or radiology finding. The term ”concerning for” can be used to indicate a doctor’s assessment of the presence of a disease based on the interpretation of imaging results. It can also be preceded by ”no” to convey the certainty of the absence of a particular disease. With the exception of the descriptive words mentioned above, all other terms listed in Fig. S1 indicate the semantic notion that the entity is absent. Appropriately, the quantity of negative descriptive words is 20×20$\times$20 × greater than that of the positive ones. Thus we propose a rebalanced factor to reweight the loss function to compensate for the dataset bias. Table 5 in the manuscript demonstrates that our re-balanced factor designed according to statistics contributes significantly to the improvements on multiple downstream tasks.
Figure S1: The statistics of descriptive words, here we only show the statistics of descriptors with a frequency greater than 10,000, otherwise, the terms are collectively included in the category labeled as ”others”.
Appendix B Selected entities
@@
Appendix C Prompt used to distill reports, and examples of distilled reports
Table S2 displays the prompt utilized for report distillation. By inputting the radiology report as the content to ChatGPT [41] with this specific prompt, we obtain a distilled report that includes more detailed semantics regarding the existence and severity of entities. Subsequently, using ChatGPT [41] to generate 10,000 distilled reports, we train a Vicuna7B [60] to further distill all the reports. Examples of these distilled reports are presented in Table S3.
Messages = [ { "role": "user", "content": f""" You are a knowledgeable radiologist. please help me analysis the medical reports and write brief and clear conclusions. Now, I will give you a medical report, as well as the mentioned entities. Please make conclusions according the medical reports, and write brief and clear conclusions as the following format: ’There is [a] [b].’ when you are sure whether the entity exists in the report, or ’There may [a] [b].’, when you are not sure whether the entity exists in the report. [a] represents adjective words describing the severeness of the entity [b]. Please generate conclusions for the entities mentioned above one by one, according to the format and do not generate other words. Please keep only one entity in a sentence, there is no need of using ’and’ or ’or’ to connect two or more words. """ }, { "role": "user", "content": f""" —Example—- Report: as compared to the lung volumes have slightly decreased. signs of mild over inflation and moderate pleural effusion persist. elongation of the descending aorta. Entities: aorta, inflation, effusion Conclusion: There is moderate pleural effusion. There is mild over inflation. There is descending aorta. —Example END—- Given the report: {report} Entities: {entity}. Conclusion: ”””}]
@@ Appendix C Prompt used to distill reports, and examples of distilled reports
Table S3: Examples of report distillation are presented in Table S3. For each report, we showcase how the semantics of our selected entities are distilled by a large language model. The distilled reports provide more specificity regarding the existence of entities.
Appendix D Detailed performances on ChestX-ray14

### 18 Mar 2024 v2

@@ 5.3 Ablation Study and Analysis
Advantages of context distillation The first row in Table 4 is trained by basic MIM and MLM. When we use the distilled easy-comprehensible report, we observe a performance improvement of 3.5% on COVIDx [50] and 0.7% on CXR14 [51]. Moreover, when we train ECAMP without the distilled report, notable performance degradation is observed, particularly on COVIDx and RSNA [43], validating the significant role of entity-aware context distillation.
Effectiveness of descriptors masking After constantly masking the critical descriptive words (as shown in Fig 2) preceding the entities, substantial performance enhancements are observed across both coarse-grained and fine-grained tasks, exceeding our initial expectations (as shown in Table 4). Accordingly, it is confirmed that descriptors masking can better leverage the supervised signals in medical reports.
Sensitive analysis of the re-balanced factor In the performance increment between the 3rd and the 4th rows as well as the 7th row and the last row of Table 4, we highlight the effectiveness of our proposed re-balanced factor. The AUC score increases by 0.9% and 0.8% respectively on CXR14, while the ACC shows sequential improvements of 0.7% and 1.5%. It is evident that the re-balanced factor primarily contributes to the performance of classification tasks. We further provide hyper-parameter analysis of $\lambda_{neg}$ corresponding to different imbalance ratios between positive and negative expressions. As shown in Table. 5, setting $\lambda_{neg}$ to $0.05$ according to the imbalance ratio (20:1) leads to best performance.
Performance of context-guided SR As listed in the 4th row and the last row of Table 4, employing context-guided SR provides performance gains from 83.0% to 86.3% on COVIDx. Notably, on SIIM, the Dice score improves from 66.6% to 67.3% and mAP sees a notable increase of 4.9% on RSNA. This underscores the efficacy of context-guided super resolution in aiding pre-trained models to capture prior knowledge of visual texture structure, particularly enhancing performance in fine-grained visual downstream tasks.
Contributions of MSCF We exclusively integrate the global vision feature $f_{v}^{g}$ and local aggregated feature $f_{a}^{l}$ to predict the masked words respectively. The results in Table 4 validate that, only fusing $f_{v}^{g}$ decreases the performances on low-level downstream tasks (segmentation and detection), while the classification performance relies on well-trained $f_{a}^{l}$. It is confirmed that MSCF successfully contributes to learning local and global representations and concurrently achieving SOTA performance on multi-scale downstream tasks.
Figure 4: Cross-attention maps between diagnosis and vision patterns learned by local aggregated features in MSCF, which successfully localizes the pathology region of various diseases.
Representation analysis As the attention map of the local vision feature and the corresponding distilled reports visualized in Fig. 4, the text modality of ECAMP captures the correct pathological regions of lung collapse (marked by red boxes) through cross-modal attention with the visual modality. Furthermore, there is no particularly disease-sensitive cross-attention for healthy images. These results support the ability of the MSCF to learn accurate and discriminative fine-grained representations, which is crucial for improving the performance of local downstream tasks.
@@ 6 Impact, Limitation, and Conclusion
Appendix
Appendix 0.A Statistics of descriptive words
As illustrated in Fig. S1, we present the statistics of our defined descriptive words across 227,835 reports. Specifically, we focus on descriptors with a frequency greater than 10,000 among the total of 31,597 terms. Notably, only "may be" and "there is" imply uncertainty about the existence of the entity, such as a disease or radiology finding. The term "concerning for" can be used to indicate a doctor’s assessment of the presence of a disease based on the interpretation of imaging results. It can also be preceded by "no" to convey the certainty of the absence of a particular disease. With the exception of the descriptive words mentioned above, all other terms listed in Fig. S1 indicate the semantic notion that the entity is absent. Appropriately, the quantity of negative descriptive words is 20×20$\times$20 × greater than that of the positive ones. Thus we propose a rebalanced factor to reweight the loss function to compensate for the dataset bias. Table 5 in the manuscript demonstrates that our re-balanced factor designed according to statistics contributes significantly to the improvements on multiple downstream tasks.
Figure S1: The statistics of descriptive words, here we only show the statistics of descriptors with a frequency greater than 10,000, otherwise, the terms are collectively included in the category labeled as "others".
Appendix 0.B Selected entities
@@
Appendix 0.C Prompt used to distill reports, and examples of distilled reports
Table S2 displays the prompt utilized for report distillation. By inputting the radiology report as the content to ChatGPT [36] with this specific prompt, we obtain a distilled report that includes more detailed semantics regarding the existence and severity of entities. Specifically, using GPT3.5-turbo API [36] setting temperature with 0 to generate 10,000 distilled reports, we train a Vicuna7B [8] to further distill all the reports. Examples of these distilled reports are presented in Table S3.
@@ Appendix 0.C Prompt used to distill reports, and examples of distilled reports
Messages = [ { "role": "user", "content": f""" You are a knowledgeable radiologist. please help me analysis the medical reports and write brief and clear conclusions. Now, I will give you a medical report, as well as the mentioned entities. Please make conclusions according the medical reports, and write brief and clear conclusions as the following format: ’There is [a] [b].’ when you are sure whether the entity exists in the report, or ’There may [a] [b].’, when you are not sure whether the entity exists in the report. [a] represents adjective words describing the severeness of the entity [b]. Please generate conclusions for the entities mentioned above one by one, according to the format and do not generate other words. Please keep only one entity in a sentence, there is no need of using ’and’ or ’or’ to connect two or more words. """ }, { "role": "user", "content": f""" —Example—- Report: as compared to the lung volumes have slightly decreased. signs of mild over inflation and moderate pleural effusion persist. elongation of the descending aorta. Entities: aorta, inflation, effusion Conclusion: There is moderate pleural effusion. There is mild over inflation. There is descending aorta. —Example END—- Given the report: {report} Entities: {entity}. Conclusion: """}]
Table S3: Examples of report distillation are presented. For each report, we showcase how the semantics of our selected entities are distilled by a large language model. The distilled reports provide more specificity regarding the existence of entities.
@@
Appendix 0.D Detailed performances on ChestX-ray14